OpenStudy (kainui):

How would I go about finding the value of this infinite sum? I just want to know the strategy, I have the answer.

OpenStudy (kainui):

\[\sum_{n=0}^{\infty} \frac{n}{10^n} \]

OpenStudy (anonymous):

humm 1/10+2/100+3/1000+.....=0.1234567910111213141516.......

OpenStudy (anonymous):

hehe :P

OpenStudy (anonymous):

how much ur gonna type :O

OpenStudy (kainui):

Yeah, well it's precisely 10/81. I derived it myself, but I really worked backwards and can't see how I'd work it without just luckily finding it the wrong way. Here's what I did: \[1^2=. \bar 9 ^2=(.9+.09+.009+...)^2=.9(.9+.09+.009+...)+.09(.9+.09+.009+...)+...\] Noticing that there will only be one .81 term, two .081 terms, etc... \[1*.81+2*.081+3*.0081+...\] Yeah you can see how I got there now, yeah?

OpenStudy (anonymous):

mmm just try to expand the sumation thats is the tric

OpenStudy (kainui):

But I want to go backwards without knowing this weird way. I don't really see how expanding the summation really works.

OpenStudy (anonymous):

it is work :) 1/10 + 2/100 +3/1000+4/10000+...=0.1+0.02+0.003+0.0004+...

OpenStudy (anonymous):

ill show u another method

OpenStudy (kainui):

Yeah but your way doesn't get the exact answer, 10/81 which is really what I want, not a decimal approximation.

OpenStudy (anonymous):

if u wanna something more neat

ganeshie8 (ganeshie8):

\(\large \sum \limits_{n=0}^{\infty}x^n = \dfrac{1}{1-x}\) differentiate both sides and plugin x = 1/10

ganeshie8 (ganeshie8):

\(1^2=. \bar 9 ^2\\=(.9+.09+.009+...)^2 \\ =.9(.9+.09+.009+...) +.09(.9+.09+.009+...)+...\)

OpenStudy (kainui):

Yeah thanks I should have tried that thanks @ganeshie8

OpenStudy (kainui):

Thanks for being helpful @BSwan I appreciate it.

ganeshie8 (ganeshie8):

hey @Kainui how did u jump from \(1*.81+2*.081+3*.0081+...\) to \(\dfrac{10}{81}\) ? im still struggling to see the link lol

OpenStudy (kainui):

Oh well from there that's really just all 1, so I made it into the summation: \[1=\sum_{n=1}^{\infty} \frac{n*.81}{10^{n-1}}\] combining the constant parts .81 and 10 turns it into: \[1=8.1\sum_{n=1}^{\infty} \frac{n}{10^n}\] so I pulled it out and got 1/8.1 which is the same as 10/81.

OpenStudy (anonymous):

so by chance u knew its 81 ?? or did u tried another method and get that??

OpenStudy (kainui):

There was no chance involved, uhh... I don't really know what part is unclear but I can try to explain anything that seems like I made a guess or something. I stayed up all night so I might have accidentally left out important details, so tell me so I can help you guys out in understanding what's going on. @_@

OpenStudy (anonymous):

Let: x= 1/10 + 2/100 + 3/1000 + ... 10x = 1 + 2/10 + 3/100 + ... Subtracting , 9x = 1 + 1/10 + 1/100 + 1/1000 + ... = 10/9 Hence: x = 10/81 Why to complicate things when it can be done simply

OpenStudy (kainui):

@No.name that works too I guess it just didn't seem very obvious to me that we could do that either. Nifty tricks all around. =P

ganeshie8 (ganeshie8):

brilliant ! thats the euler approach :P i think i got Kainui's method... its not complicated - its just a simple matter of collecting terms in a product : (a+b+c+...)(x+y+z+...)

OpenStudy (anonymous):

Euler again , how nice is it :P

OpenStudy (anonymous):

I don't know euler approach but that's common sense

ganeshie8 (ganeshie8):

its gauss i guess lol^

OpenStudy (kainui):

Now you've got me thinking I can work the approach backwards to get even messier ones to behave like a geometric series 2 steps away rather than 1 step away. Now I want to generalize it to n steps away and play around lol.

OpenStudy (anonymous):

he was 9 yrs old i guess :P

OpenStudy (anonymous):

who

OpenStudy (anonymous):

gauss

OpenStudy (anonymous):

what r u talking abt gauss 9 year old?

OpenStudy (anonymous):

-.- nvm

ganeshie8 (ganeshie8):

lol me neither, i meant gauss would have approached this problem like the way @No.name solved it :)

OpenStudy (anonymous):

What is Bswan talking about ?

OpenStudy (anonymous):

i said i cant see gauess or eluer in this lol when ganeshie said about gauss and provide the link i said he was 9 thats all ^^

ganeshie8 (ganeshie8):

She is saying, Gauss was 9 years old when he figured out the formula for sum of first 100 numbers

OpenStudy (anonymous):

oh lol okay

OpenStudy (kainui):

I found an interesting series that you can solve because of this: \[\sum_{n=1}^{\infty} \frac{n^2}{10^n}\] Kinda interesting what kinds of things you can play around with, anyone know any other interesting series?

OpenStudy (anonymous):